# No Filter: Cultural and Socioeconomic Diversity in Contrastive Vision–Language Models

### v2

@@ 1 Introduction
In line with [2], we focus on contrastive learning for several reasons. These models have a wide range of applications, e.g. zero-shot classification and cross-modal retrieval, and are being increasingly adopted in critical domains like healthcare [75, 51], and as a backbone for other models [72, 71, 40]. We study SigLIP [74], the currently best-performing and a widely used CLIP-style model.
@@ 2 Preliminaries
where $t>0$ and $b$ are learnable temperature and bias parameters.We follow Zhai et al., [74] in most aspects: we use a Vision Transformer (ViT) [18] for images and a Transformer [59] for text, both of size Base (B), with an embedding dimension of $768$. Images are resized to $256\times 256$ resolution and we use $16\times 16$ patch size. Text input is tokenized using a model trained on mC4 [68], a Common Crawl-based dataset covering 101 languages, with a vocabulary size of 250k. We keep a maximum of $64$ tokens. We use the modified Adafactor optimizer [53, 73] for all our experiments with an initial learning rate of $10^{-3}$, reverse square root decay, weight decay of $10^{-4}$ and 50k warmup and cooldown steps. Training batch size is 16K. Models are developed in the Big Vision codebase [6] using Tensor Processing Units (TPUs) [31]. Each model is trained on 10 B image–text pairs (roughly 610 k steps) for approximately 40 k TPUv2 core-hours.
@@ 3.1 No filter for improved cultural diversity
Image $\rightarrow$ Text 50.60 ±$\pm 1.54\%$ 49.10 ±$\pm 0.28\%$ 49.74 ±$\pm 1.03$ $-0.86\%$
@@
Image $\rightarrow$ Text 62.73 ±$\pm 1.28\%$ 60.70 ±$\pm 1.38\%$ 62.02 ±$\pm 1.41$ $-0.71\%$
@@
Indeed, a more fine-grained analysis of zero-shot classification on Dollar Street confirms that filtering to English-only training data disproportionately hurts low-income and non-Western communities, see Figure 1 (right) and Figure 3 (left). In MaRVL, performance for all regions (except TA) tend to benefit from using globally diverse training data, as shown in Figure 3 (right). The difference in classification accuracy for different income groups and geographic regions is a worrying signal of the inherent biases of filtering to English-only data. Removing the English language filter unsurprisingly leads to a drop in performance for images from Western countries, as exemplified by ImageNet and COCO benchmarks, but it significantly improves performance for the rest of the world.
@@ 3.2 Few-shot geo-localization
Figure 3: Filtering to English-only data further exacerbates existing socioeconomic disparities. left: Zero-shot classification results for Dollar Street, disaggregated by income level ($x$-axis). The performance difference between en and globe-tl is larger for lower-income households. right: MaRVL Concepts classification accuracy disaggregated by each of the five languages/regions: Pretraining on globe-tl improves performance for Indonesian, Turkish and Mandarin Chinese and yields a similar performance to en for Tamil and Swahili.
@@ 4 Related Work
A range of prior work has studied biases in zero-shot classifiers related to a range of sensitive attributes including gender, race and age [1, 24, 26, 12] and found that CLIP models perpetuate biases present in the training data [7, 23, 2]. More recent work extends this analysis to zero-shot performance across groups of different income levels and geographic regions [42, 77]. While several of these papers highlight the central role of training data and the potentially large impact of design choices such as the data source or filtering techniques [22] or translation of English captions for cross-modal multilingual encoders [10, 11, 44], to our knowledge, we are the first to study the impact of image–text pair filtering and text translation on cultural understanding in contrastive VLMs.
@@
The closest work to ours is [47], which argues that progress in global data (Dollar Street and GeoDE) has been much slower than the progress on ImageNet. Unlike their work, however, we study the impact of the training data mixture, study the impact of translation, suggest a setup where both types of metrics can be improved, and propose geo-localization as an evaluation metric. We also consider a broader set of datasets in our study, such as XM-3600, GLDv2, and MaRVL.

### v3

@@ 1 Introduction
In line with Alabdulmohsin et al., [2], we focus on contrastive learning for several reasons. These models have a wide range of applications, e.g. zero-shot classification and cross-modal retrieval, and are being increasingly adopted in critical domains like healthcare [76, 52], and as a backbone for other models [73, 72, 41]. We study SigLIP [75], the currently best-performing and a widely used CLIP-style model. SigLIP and CLIP operate on the same principle of aligning representations/embeddings for texts and images and the difference is only in the choice of the loss function.
@@ 2 Preliminaries
where $t>0$ and $b$ are learnable temperature and bias parameters.We follow Zhai et al., [75] in most aspects: we use a Vision Transformer (ViT) [18] for images and a Transformer [60] for text, both of size Base (B), with an embedding dimension of $768$. Images are resized to $256\times 256$ resolution and we use $16\times 16$ patch size. Text input is tokenized using a model trained on mC4 [69], a Common Crawl-based dataset covering 101 languages, with a vocabulary size of 250 k. We keep a maximum of $64$ tokens. We use the modified Adafactor optimizer [54, 74] for all our experiments with an initial learning rate of $10^{-3}$, reverse square root decay, weight decay of $10^{-4}$ and 50 k warmup and cooldown steps. Training batch size is 16 k. Models are developed in the Big Vision codebase [6] using Tensor Processing Units (TPUs) [32]. Each model is trained on 10 B image–text pairs (roughly 610 k steps) for about 40 k TPUv2 core-hours, so they are compared on a compute-matched regime.
@@ 3.1 No filter for improved cultural diversity
Image $\rightarrow$ Text 50.60 ±$\pm 1.54\%$ 49.10 ±$\pm 0.28\%$ 49.74 ±$\pm 1.03\%$ $-0.86\%$
@@
Image $\rightarrow$ Text 62.73 ±$\pm 1.28\%$ 60.70 ±$\pm 1.38\%$ 62.02 ±$\pm 1.41\%$ $-0.71\%$
@@
Indeed, a more fine-grained analysis of zero-shot classification on Dollar Street confirms that filtering to English-only training data disproportionately hurts low-income and non-Western communities; see Figure 1 (right) and Figure 3 (left). In MaRVL, performance for all regions (except TA) tend to benefit from using globally diverse training data, as shown in Figure 3 (right). The difference in classification accuracy for different income groups and geographic regions is a worrying signal of the inherent biases of filtering to English-only data. Removing the English language filter unsurprisingly leads to a drop in performance for images from Western countries, as exemplified by ImageNet and COCO benchmarks, but it significantly improves performance for the rest of the world.
@@ 3.2 Few-shot geo-localization
Figure 3: Filtering to English-only data further exacerbates existing performance disparities across socioeconomic subgroups. left: Zero-shot classification results for Dollar Street, disaggregated by income level ($x$-axis). The performance difference between en and globe-tl is larger for lower-income households. Also, the performance disparity between the lowest and highest income groups is 32.5% in en (from 29.9% in \$0-200 income group to 62.4% in \$1998+ income group), but this gap is reduced (improved) to 27.4% in globe-tl. right: MaRVL Concepts classification accuracy disaggregated by each of the five languages/regions: Pretraining on globe-tl improves performance for Indonesian, Turkish and Mandarin Chinese and yields a similar performance to en for Tamil and Swahili.
@@ 3.4 Bridging the gap
3.5 Quality Filters
We apply quality filtering using an internal model trained on image-text pairs to calculate the image-text similarity score, in order to assess if our empirical findings continue to hold in this context. The model was trained on global data to make sure it accurately assesses quality for global data, of which English data is a substantial fraction, and its threshold was tuned to balance quality and quantity. We filter out about 60% of the data in our experiments. Appendix B shows that our main findings continue to hold even when quality filters are applied. For instance, quality-filtered globe-tl performs better than quality-filtered en on 0-shot Dollar Street, GLDv2, and GeoDE but performs worse on Western-oriented benchmarks, such as ImageNet and COCO retrieval. In addition, the improvement in quality-filtered globe-tl over quality-filtered en is particularly significant for few-shot geo-localization tasks.
@@ 4 Related Work
A range of prior work has studied biases in zero-shot classifiers related to a range of sensitive attributes including gender, race and age [1, 25, 27, 12, 24] and found that CLIP models perpetuate biases present in the training data [7, 23, 2]. More recent work extends this analysis to zero-shot performance across groups of different income levels and geographic regions [43, 78]. While several of these papers highlight the central role of training data and the potentially large impact of design choices such as the data source or filtering techniques [22] or translation of English captions for cross-modal multilingual encoders [10, 11, 45], to our knowledge, we are the first to study the impact of image–text pair filtering and text translation on cultural understanding in contrastive VLMs.
@@
The closest work to ours is [48], which argues that progress in global data (Dollar Street and GeoDE) has been much slower than the progress on ImageNet. Unlike their work, however, we study the impact of the training data mixture, study the impact of translation, suggest a setup where both types of metrics can be improved, and propose geo-localization as an evaluation metric. We also consider a broader set of datasets in our study, such as XM3600, GLDv2, and MaRVL.
@@ A.1 Model Card
Appendix B Impact of Quality Filters
To assess the impact of quality filters on our findings, we train two models, en and globe-tl, on 1B image-text pairs, following the same training setup used in the paper. Table 4 shows a summary of these results. We observe that our empirical findings also hold in this setting. For instance, quality-filtered globe-tl performs better than quality-filtered en on 0-shot Dollar Street, GLDv2, and GeoDE but performs worse on Western-oriented benchmarks, such as ImageNet and COCO retrieval. In addition, the improvement in few-shot geo-localization for quality-filtered globe-tl over quality-filtered en is particularly significant.
Table 4: Applying quality filters to SigLIPs does not change the primary conclusions. Filtering training data to English image-text pairs continues to negatively impact cultural diversity even though it improves performance on standard benchmarks.
en globe-tl en vs. globe-tl
Culturally diverse zero-shot evaluations
Dollar Street $46.05\%$ $48.28\%$ $+2.23\%$
GLDv2 $28.21\%$ $30.67\%$ $+2.46\%$
GeoDE $90.53\%$ $90.57\%$ $+0.04\%$
Prevalent Western-oriented benchmarks
0-shot ImageNet $66.96\%$ $66.32\%$ $-0.64\%$
[…]
10-shot geo-localization
Dollar Street (country) $9.40\%$ $9.82\%$ $+0.42\%$
GeoDE (country) $10.10\%$ $14.73\%$ $+4.63\%$
GeoDE (region) $21.97\%$ $28.22\%$ $+6.25\%$